### v1

# Continued radio observations of the persistent radio source associated with FRB20190520B provides insights into its origin

Abstract
Follow-up studies of persistent emission from Fast Radio Burst (FRB) sources are critical for understanding their elusive emission mechanisms and the nature of their progenitors. This work presents new observations of the persistent radio source (PRS) associated with FRB20190520B. We observe a gradual decay in the PRS brightness, which is punctuated by periods of brightening and dimming at both 1.5 GHz and 3 GHz. Furthermore, our low-frequency ($<1$ GHz) observations—the first for this source—reveal evidence of a spectral break which can be attributed to absorption processes. Interpreted within the framework of the magnetar wind nebula model, our data constrain the age of the magnetar progenitor to $53^{+17}_{-10}$ years, broadly consistent with previous work. Assuming the observed 1.5 GHz variability is driven by scintillation, we derive a conservative lower limit for the source’s radius of $>0.52$ pc, for a screen at 10 kpc. The observations presented here challenge the predictions of the previously published best-fit hypernebula model for this source.
Radio transient sources (2008), Transient sources (1851), Radio sources (1358), Radio bursts (1339)
software: CASA (CASA Team et al., 2022), CASA CAPTURE (Kale & Ishwara-Chandra, 2021), emcee (Foreman-Mackey et al., 2013), ChainConsumer (Hinton, 2016), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), NumPy (Harris et al., 2020), SciPy (Virtanen et al., 2020), Matplotlib (Hunter, 2007)
1 Introduction
Fast Radio Bursts (FRBs) are highly energetic $\sim$ millisecond duration radio transients (Lorimer et al., 2007). The observations of FRBs using various radio facilities have revealed that some FRBs repeat, while others appear to be non-repeating (Petroff et al., 2022). Among the FRBs that repeat, only two have shown evidence of a periodicity in their activity windows (Petroff et al., 2022, and refs. therein). FRBs experience propagation effects due to interaction and scattering before reaching our detectors. Some propagation effects include: dispersion measure (DM)—the delay in arrival of low frequency radiation compared to the high frequency radiation due to interaction with ionized interstellar medium; rotation measure (RM)—Faraday rotation of the polarization angle due to magnetic field along the line of sight. The measured DM of the FRB can be expressed as a sum of the DM contribution from the Milky Way, the interstellar medium, the intergalactic medium, and the local DM from the host galaxy of the FRB.
The origin of FRBs is still an open question, with many proposed models (Platts et al., 2019). The discovery of an FRB-like burst from galactic magnetar SGR1935+2154 has bolstered interest in magnetar-based models as the leading candidates for explaining FRBs (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). To date, a total of 865 FRBs have been identified, of which 71 are repeaters (Xu et al., 2023). 96 FRBs have been localized to their host galaxies, providing rich insights into FRB environments (see e.g. Chatterjee et al., 2017; Tendulkar et al., 2021; Marcote et al., 2020; Niu et al., 2022).
Among the localized repeaters, two FRBs—FRB20121102 (Chatterjee et al., 2017) and FRB20190520B (Niu et al., 2022)—are of particular interest given that a compact persistent radio source (PRS) have been observed associated with both, making them key candidates to test FRB progenitor models. A possible PRS has also been reported for FRB 20201124A by Bruni et al. (2024).
FRB20190520B was discovered by Five-hundred-meter Aperture Spherical radio Telescope (FAST) on May 20, 2019 as part of the CRAFT Survey (Li et al., 2018). It has been localized to a star forming dwarf galaxy at a redshift of $z=0.241$ (Niu et al., 2022; Chen et al., 2025). The average DM of bursts from this source is 1207 cm-3pc. Initially, the DM contribution from the host was thought to be exceptionally large $\sim 900$ cm-3pc. However, Lee et al. (2023) investigated the effect of foreground galaxies and estimated a substantially lower host DM. (also see Bhardwaj et al., 2025b). Observations of bursts from this FRB using the Green Bank Telescope (GBT) and Parkes telescope by (Anna-Thomas et al., 2023) have shown that the RM changes sign twice, pointing to a change in the magnetic field component parallel to the propagation, possibly due to a turbulent magnetized screen surrounding the FRB source. Bhandari et al. (2023) used Very-long-baseline interferometry (VLBI) observations to constrain the size of the PRS $<9$ pc and colocated it to within $80$ pc of the FRB. This PRS exhibits a flat spectrum (slope of $-0.4$ at 2020 and $-0.33$ at 2021 epochs) above 1 GHz accompanied by a $\sim 20$% decline in broadband flux between the 2020 and 2021 epochs (Zhang et al., 2023). The PRS has not been detected at frequencies $<1$ GHz in the past.
Some leading models proposed to explain persistent radio emission from FRB sources are— Magnetar Wind Nebula (MWN) model (Margalit & Metzger, 2018) and hypernebula model (Sridhar & Metzger, 2022). The Magnetar Wind Nebula model attributes persistent radio emission to synchrotron emission from magnetized electron-ion interactions with the nebula, powered by flaring activity of a young magnetar. The Hypernebula model explains the persistent emission as the interaction of winds emitted by an accreting black hole with the surrounding medium. Details of the interpretations of these models are discussed in Section 3.
In this paper, we present the first ever low frequency observations of the persistent radio source associated with FRB20190520B. We interpret the temporal and spectral evolution of the persistent emission by combining our new observations and previous observations of the source. Section 2 describes the observations and data reduction performed, followed by Section 3 covering the analysis of the observations. Finally, we conclude with a discussion of our analysis in Section 4.
2 Observations
2.1 uGMRT observations and archival data
We observed the PRS associated with FRB20190520B using the wideband receiver backend of upgraded Giant Metrewave Radio Telescope (uGMRT) in three frequency bands: band 3 (central frequency, $\nu_{c}$ = 400 MHz, bandwidth, BW = 200 MHz), band 4 ($\nu_{c}$ = 750  MHz, BW = 400 MHz) and band 5 ($\nu_{c}$ = 1260 MHz, BW = 400 MHz) between 2023 June 16 and 2023 June 20 (Proposal 44_039, PI: Balasubramanian). The band 3 observations under this project were split over two days. We also obtained three epochs of observations each in band 4 and band 5 between 2024 September 05 and 2024 September 21 (Proposal 46_126, PI: Balasubramanian). In addition to our observations, we included archival uGMRT data from proposal 43_054 (PI: Yi Feng) in our analysis. Raw data were downloaded in the FITS format and converted to the CASA (CASA Team et al., 2022) measurement set format. The data were then calibrated and imaged using the automated continuum imaging pipeline CASA-CAPTURE (Kale & Ishwara-Chandra, 2021). For the band 3 data, the two measurement sets were passed onto the automated pipeline separately and then imaged after combining the calibrated measurement sets. Each pipeline run included eight rounds of self-calibration. The band 5 flux density of the persistent source was calculated using the CASA task imfit within a small circular region (of radius 2$\sim$ 2$\times$ 2 × the size of the synthesized beam at band 5) centred at the PRS coordinates. An additional 5% flux density error was added in quadrature to the error obtained from imfit to account for flux calibration errors (see Table 1). No detections were made in Band 3 or in any epoch of Band 4 observations. The upper limit values listed in Table 1 are the 3×3$\times$3 ×RMS value within a large circular region (of radius 20$\sim$ 20$\times$ 20 × the size of the synthesised beam at the respective band) centred at the position of the PRS.
2.2 VLA observations and archival data
We conducted a single L-band ($\sim$1.5/,GHz) observation of the PRS using the Jansky Very Large Array (VLA) on 2024 August 08. In addition, VLA S-band ($\sim$3 GHz) data of the PRS20190520B field were obtained from the archival data listed under proposal 23A-010 (PI: Yi Feng, taken between 2023 June 07 and 2023 June 24). We obtained the calibrated measurement set when available or started with the raw data. Raw data were calibrated using the automated VLA calibration pipeline  [^1] https://science.nrao.edu/facilities/vla/data-processing/pipeline/CIPL_654. Following calibration, the data were imaged using the automated imaging pipeline  [^2] https://science.nrao.edu/facilities/vla/data-processing/pipeline/VIPL and refined through self-calibration. The best self-calibrated image was selected, and the CASA task imfit was used in a small region around the PRS to estimate the source flux density. An additional 5% flux uncertainty was added in quadrature to account for flux calibration errors.
A summary of all measurements is listed in Table 1. Figure 1 shows the temporal variation of the PRS flux density at different frequency bands, and Figure 2 displays the variation of the spectrum over the epochs. We observe flaring and dimming episodes of the source in the $\approx 1.5$ GHz data, especially in the late 2020 epoch. The possible reason for this observation is discussed in the Section 3.2. For some observations, the image was heavily affected by radio frequency interference (RFI), resulting in an image that was insufficiently clean to use for our analysis (see Table 1 footnote).
[…]
Table 1: Summary of radio continuum observations of the persistent radio source associated with FRB20190520B. For non-detections, 3 $\sigma$ upper limits are reported, where $\sigma$ is the RMS flux density measured in a large region of the residual image. Observation dates (in two formats), frequency bands, flux densities and their associated errors, and proposal identifiers/paper references are listed.
Date Time (MJD) Frequency (GHz) Flux density (uJy) Flux density error (uJy) Reference
[…]
2020 Jul 23 59053.08 1.5 273 37
2020 Aug 18 59079.019 5.5 145 17
2020 Aug 18 59079.986 5.5 164 19
2020 Aug 29 59090.956 5.5 158 17
2020 Aug 30 59091.953 3.0 195 24
2020 Sep 12 59104.881 3.0 160 21
2020 Sep 12 59104.923 5.5 151 17
2020 Sep 13 59105.991 3.0 186 24
2020 Sep 15 59107.93 5.5 153 17
2020 Sep 19 59111.13 3.0 176 25
2020 Nov 08 59161.691 5.5 139 20
2020 Nov 14 59167.652 3.0 233 29
2020 Nov 16 59169.655 3.0 211 25
2021 Oct 01 59488.879 10.0 115 24
2021 Oct 01 59488.883 5.5 114 28
2021 Oct 01 59488.887 3.0 112 34
2021 Oct 01 59488.895 1.5 240 70
2021 Nov 07 59525.861 10.0 81 18
2021 Nov 07 59525.865 5.5 139 33
2021 Nov 07 59525.869 3.0 111 33
2021 Nov 07 59525.877 1.5 212 61
[…]
2022 Feb 27 59637.208 1.7 210 34
2022 Nov 11 59894.205 1.3 $<$194 GMRT 43_054 (This work)
2022 Nov 29 59912.125 1.3 268 44 GMRT 43_054 (This work)
2022 Dec 27 59940.131 1.3 156 26 GMRT 43_054 (This work)
2023 Jan 24 59968.036 1.3 281 41 GMRT 43_054 (This work)
2023 Feb 21 59996.167 1.3 $<$122 GMRT 43_054 (This work)
2023 Mar 21 60024.962 1.3 $<$144 GMRT 43_054 (This work)
2023 Jun 07 60102.323 3.0 157 9 VLA 23A-010 (This work)
2023 Jun 15 60110.234 3.0 173 11 VLA 23A-010 (This work)
2023 Jun 16 60111.84 0.7 $<$165 GMRT 44_039 (This work)
2023 Jun 17 60112.192 3.0 171 12 VLA 23A-010 (This work)
2023 Jun 18 60113.178 3.0 166 10 VLA 23A-010 (This work)
2023 Jun 18 60113.219 3.0 165 11 VLA 23A-010 (This work)
2023 Jun 19 60114.196 0.3 $<$195 GMRT 44_039 (This work)
2023 Jun 20 60115.172 3.0 152 10 VLA 23A-010 (This work)
2023 Jun 20 60115.214 3.0 151 11 VLA 23A-010 (This work)
2023 Jun 20 60115.255 3.0 143 10 VLA 23A-010 (This work)
2023 Jun 20 60115.297 3.0 152 10 VLA 23A-010 (This work)
2023 Jun 20 60115.584 1.3 179 18 GMRT 44_039 (This work)
2023 Jun 23 60118.262 3.0 147 13 VLA 23A-010 (This work)
2023 Jun 24 60119.168 3.0 155 15 VLA 23A-010 (This work)
2023 Jun 24 60119.209 3.0 144 13 VLA 23A-010 (This work)
2024 Aug 08 60530.985 1.5 221 13 VLA 24A-409 (This work)
2024 Sep 03 60556.588 1.3 230 24 GMRT 46_126 (This work)
2024 Sep 05 60558.585 0.7 $<$155 GMRT 46_126 (This work)
2024 Sep 12 60565.589 1.3 184 16 GMRT 46_126 (This work)
$\dagger$\$$\dagger$\$footnotemark: 2024 Sep 15 60568.594 0.7 GMRT 46_126 (This work)
2024 Sep 21 60574.329 1.3 190 20 GMRT 46_126 (This work)
$\dagger$\$$\dagger$\$footnotemark: 2024 Sep 21 60574.0 0.7 GMRT 46_126 (This work)
footnotetext: Data affected by severe radio frequency interference. Sufficiently clean image was not obtained.
Figure 1: Light curve of the PRS associated with FRB20190520B. The temporal evolution of the radio emission hints at a steady decay of the PRS flux density, with some flaring and dimming seen superimposed on the decay. Previous data plotted here are from Niu et al. (2022) and Zhang et al. (2023).
Figure 2: Spectrum of the PRS associated with FRB20190520B. The upper limits at frequencies below 1 GHz from this work shows the possibility of a spectral break. Previous data plotted here are from Niu et al. (2022) and Zhang et al. (2023).
3 Physical Interpretations
Equipped, With the newly observed multi-frequency dataset of the PRS associated with FRB20190520B, we present constraints on the physical properties of the possible progenitor of the persistent emission.
3.1 Pulsar Wind Nebula Model
Margalit & Metzger (2018) showed that a single expanding magnetized electron-ion nebula powered by a young magnetar can interact with the surrounding medium to produce synchrotron emission. The magnetar is assumed to inject energy into the nebula at a rate given by
$\dot{E}=(\alpha-1)\frac{E_{B_{\star}}}{t_{0}}\left(\frac{t}{t_{0}}\right)^{-\alpha}\hskip 5.0pt\rm{for}\hskip 5.0ptt\geqslant t_{0},\hskip 5.0pt\alpha>1$ (1)
where $E_{B_{\star}}$ is the free magnetic energy of the magnetar, $\alpha$ is a powerlaw index, and $t_{0}$ is the onset of the active period of the magnetar. The energy injection drives electrons through the surrounding medium, producing the observed persistent emission. At frequencies $\nu$ above the characteristic synchrotron self-absorption frequency $\nu_{SSA}$, the synchrotron luminosity decays as follows (see Margalit & Metzger, 2018, for detailed calculations)
$L_{\nu}\propto\nu^{-\left(\frac{\alpha-1}{2}\right)}t^{-\frac{\alpha^{2}+7\alpha-2}{4}}$ (2)
We adopt Equation 2 and modify it as follows
$F(\nu,t)=A\,\nu^{-\left(\frac{\alpha-1}{2}\right)}\left(\frac{t_{\rm{obs}}+t_{\rm{age}}}{t_{\rm{age}}}\right)^{-\left(\frac{\alpha^{2}+7\alpha-2}{4}\right)},$ (3)
where $A$ is a scaling constant, $t_{\rm{obs}}$ is the time since the first observation of the PRS, $t_{\rm{age}}$ is the age of the persistent source at the time of the first observation. We perform an MCMC fit of Equation 3 using the emcee package (Foreman-Mackey et al., 2013), incorporating all the detections listed in Table 1. The posterior distributions were analysed using ChainConsumer (Hinton, 2016) resulting in the best-fit parameters listed in Table 2. Figure 3 shows the best fit light curves for different frequencies, following the color-code of the data points. The corresponding best fit spectra are shown in Figure 4. The best-fit age of the persistent source, $t_{\rm{age}}$, is $53^{+17}_{-10}$ years, consistent with the allowed age limits derived in Bhandari et al. (2023) and slightly higher than the 16–22 year estimate from Zhao & Wang (2021). The best-fit energy injection power law index is $\alpha=1.75^{+0.09}_{-0.08}$. Excluding the flaring and dimming events observed in late 2022 does not significantly alter the best-fit parameter values.
Figure 3: Light curve of the FRB20190520B PRS, showing the joint spectro-temporal fit using Equation 3. This fit gives an estimate of the age of the persistent source to be $t_{\rm{age}}=53$ years.
Figure 4: Spectrum of the PRS of FRB20190520B showing the joint spectro-temporal fit using Equation 3. The color of the dashed lines represent the different epochs shown in the same colors. The shaded blue region shows the uGMRT upper limits that suggest a break in the spectrum at frequencies $<1\,GHz$.
Table 2: MCMC posterior best fit parameters for the decay using Equation 3.
Parameter Prior type Bounds Posterior value
$\alpha$ uniform (1, $\infty$) $1.75^{+0.09}_{-0.08}$
$t_{\rm{age}}$ (yr) uniform (4,1900) $53^{+17}_{-10}$
$A$ (uJy) uniform (1, $\infty$) $280^{+20}_{-18}$
In addition to the flux decay observed at frequencies $\nu>1$ GHz, the upper limits from our uGMRT observations at $\nu<1$ GHz (from the 2023 and 2024 epochs) suggest a spectral break between 700 MHz and 1.3 GHz (see blue shaded portion in Figure 4). At the observed timescales, such a spectral break is consistent with synchrotron self-absorption. However, the available upper limits are insufficient to constrain the synchrotron self-absorption frequency. Deeper observations at sub-GHz frequencies will be required to quantitatively determine the location of the spectral break.
3.2 Scintillation as a source of PRS variability
Zhang et al. (2023) analysed temporal variability above 1 GHz and found no significant variation in most bands. A marginal variability was seen at 3 GHz. We extend the same variability analysis, adding our 1.3 GHz observations to the 1.5 GHz observations reported by Zhang et al. (2023). One of the causes of variability of a compact radio source is interstellar scintillation. Scintillation arises due to inhomogeneities in the ionized ISM and can cause apparent flux variations in compact radio sources. We follow the formalism introduced in Narayan (1992) and further explored in Walker (1998) for our analysis here. The scattering strength parameter $\xi$ describes the scintillation due to a screen at a distance $D$ from the observer. $\xi=1$ corresponds to the critical value at which phase changes are substantial, across the characteristic first Fresnel zone $\theta_{F}=\sqrt{c/2\pi\nu D}$, where $\nu$ is the frequency of observation. Now, $\xi$ is expressed as $(\nu_{0}/\nu)^{17/10}$, where $\nu_{0}$ is the transition frequency between the weak scattering ($\xi\ll 1$ and $\nu_{0}<\nu$), and strong scattering ($\xi\gg 1$ and $\nu_{0}>\nu$) regimes. This transitional frequency has been estimated to be $\nu_{0}=12.53$ GHz using the galactic free electron density model in pyNE2001 (Cordes & Lazio, 2002). Given $\nu_{0}=12.53$ GHz, we are in the strong scattering regime as $\nu_{0}>\nu$ (as $\nu=1.5\,GHz$). There are two possible scintillation scenarios: diffractive scintillation and refractive scintillation. Diffractive scintillation is relevant at characteristic frequency bandwidth, of $\xi^{-2}\nu=\nu(\nu_{0}/\nu)^{17/5}$ which is $\sim 1$ MHz for our observations. Our observations span a larger bandwidth, washing out diffractive scintillation. Hence, refractive scintillation is the relevant mechanism here. For, refractive scintillation, the expected modulation index is
$m_{\rm{exp}}=\xi^{-1/3}=\left(\frac{\nu}{\nu_{0}}\right)^{17/30}=0.29$ (4)
Assuming a scattering disk at a distance of $D=1$ kpc from us, the size of the Fresnel zone at $\nu=1.5$ GHz is $\theta_{F}=\sqrt{c/2\pi\nu D}=6.74$. Therefore, the angular size of the screen is
$\theta_{r}=\theta_{F}\xi=276.6\,\rm{mas}$ (5)
This angular size of the screen can be converted to a projected radius of the screen at the distance of the source using the luminosity distance of the host galaxy (i.e. $d_{\rm{lum}}=1218$ Mpc) as follows:
$R_{r}=\theta_{r}d_{\rm{lum}}/2=5.0\times 10^{18}\,\rm{cm}=1.6\,\rm{pc}$ (6)
The characteristic timescale associated with the refractive scintillation from this screen is
$t_{r}=2\left(\frac{\nu_{0}}{\nu}\right)^{11/5}=10.2\,\rm{days}$ (7)
Given this expected timescale, we compute the modulation index of the observed data in the frequency range 1.3–1.7 GHz from Table 1, by averaging consecutive data that are within the refractive scintillation timescale, $t_{r}$, using the following expression:
$m_{\rm{obs}}=\frac{1}{\overline{F_{i}}}\sqrt{\frac{N}{N-1}\left(\overline{F_{i}^{2}}-\overline{F_{i}}^{2}\right)}$ (8)
where $F_{i}$ are the flux density measurements, $\overline{F_{i}}$ is the mean of the flux density measurements, $\overline{F_{i}^{2}}$ is the mean of the squares of $F_{i}$ and $N$ is the total number of observations.
We obtain the observed modulation index $\bm{m_{\rm{obs}}=0.18}$. The observed modulation of the PRS flux is less than that expected from this region of the sky, i.e. $\bm{m_{\rm{obs}}<m_{\rm{exp}}}$. This observation can be interpreted in the following ways:
[…]
  • All the variability observed is due to scintillation, but the pyNE2001 model estimate of the transition frequency is not accurate. If this is the case, we can estimate the transition frequency required to produce the observed modulation. Table 3 lists the expected transition frequency, $\nu_{0}$ which can produce the observed modulation index at $\nu=1.5\,$GHz.
  • All the variability observed is intrinsic to the source. This implies that the size of the source $\theta_{s}$ is larger than the size of the screen $\theta_{r}$ and therefore, refractive scintillation (which assumes a point source) is not relevant here. The projected size of the screen $R_{r}$ tabulated in Table 3 can be used as a conservative lower limit of the size of the source in this case.
At this point, we also note that there is a possibility of both intrinsic variability and scintillation to contribute to the observed variability. It is not trivial to separate the two without independent constraints on the size of the source and the properties of the scattering screen.
For completeness, we also repeat this analysis after removing the expected decay from the pulsar wind nebula model (Section 3.1). The inferred size and timescale constraints remain broadly consistent in both cases.
Figure 5: Constraint on the radius of the PRS source using the projected size of the screen $R_{r}$. The figure is adapted from Bhandari et al. (2023), after correcting for the radius by dividing all limits by 2.
Table 3: Scintillation analysis and predictions at 1.5 GHz
Dataset $t_{r}$ (days) $m_{\rm{exp}}$ $m_{\rm{obs}}$ Characteristic sizes modified $t_{r}$ [1em] (days) modified $\nu_{0}$ [1em] ( GHz)
$D=0.1$ kpc $D=1$ kpc $D=10$ kpc
$\theta_{F}$ ($\mu$as) $R_{r}$ (pc) $R_{s}$ (pc) $\theta_{F}$ ($\mu$as) $R_{r}$ (pc) $R_{s}$ (pc) $\theta_{F}$ ($\mu$as) $R_{r}$ (pc) $R_{s}$ (pc)
Data as it is 10.20 0.29 0.18 21.31 5.17 7.78 6.74 1.63 2.46 2.13 0.52 0.78 15.37 29.12
Removed best fit decay 0.16 7.36 2.33 0.74 14.54 37.59
3.3 Hypernebula Model
[…]
3.4 AGN or IMBH Origin for the Persistent Radio Source
An alternative scenario is that the persistent radio source (PRS) associated with FRB20190520B is physically unrelated to the FRB engine, and instead arises from accretion onto an massive black hole as discussed by Anna-Thomas et al. (2023). In this picture, the PRS represents compact synchrotron emission from a radio-loud, low-Eddington active galactic nucleus (AGN), while the FRB originates from a separate compact object—such as a magnetar—embedded within the same nuclear environment.
Recent work by Dong et al. (2024) has systematically compared the radio spectra of persistent sources in nearby dwarf galaxies—including known low-mass AGNs—with those of FRB-associated PRSs, including FRB20190520B and FRB 20121102A. They find that several AGNs in dwarf hosts exhibit flat or slightly inverted radio spectra and luminosities consistent with both FRB PRSs, suggesting that an AGN origin cannot be ruled out based on spectral properties alone.
[…]
A similar interpretation has been proposed for the PRS associated with FRB 20121102A. Bhardwaj et al. (2025a) argued that its long-term flux stability, compact size ($<0.35$ pc), and placement on the radio-loud fundamental plane of black hole activity can be explained by a low-Eddington AGN powered by an IMBH. Given the morphological and spectral similarities between the two sources, a comparable AGN origin for the PRS of FRB20190520B remains observationally viable. However, as with FRB 20121102A, this model requires that the FRB bursts themselves arise from a distinct, co-located source within the broader accretion environment.
Further high-resolution radio monitoring and deep X-ray observations will be essential to distinguish between AGN and magnetar wind nebula interpretations and to assess whether FRB sources and persistent emitters are causally connected or merely cohabit similar extreme environments.
4 Conclusion
In this paper, we present broadband radio observations (0.3–-10 GHz) of the persistent radio source (PRS) associated with FRB20190520B. We analyze the data under different proposed models and comment on the physics responsible for the observed emission. The findings are summarized below.
  • We observe a slow decay in the brightness of the PRS with time across all frequency bands. There are episodes of brightening and dimming seen in the 1.5 GHz and 3 GHz data.
  • The first low frequency ($<1\,$GHz) observations of this PRS hint at a spectral break between $\approx 0.7$ GHz and $\approx 1.3$ GHz, which can be attributed to synchrotron self-absorption. It is not possible to estimate the self-absorption frequency using only the upper limits from the low-frequency observations presented here.
  • We fit the decaying light curves at all frequencies using the decay equation described in Equation 3, derived from the magnetar wind nebula model. This resulted in an estimate of the powerlaw index for energy injection $\alpha=1.75^{+0.09}_{-0.08}$ and the age of the magnetar is $t_{\rm{age}}=53^{+17}_{-10}$. These values are in agreement with constraints from previous observations of the PRS (Bhandari et al., 2023; Zhao & Wang, 2021).
  • We also comment on the contribution of scintillation to the variability of the PRS at 1.5 GHz. Through this analysis, a conservative lower limit on the size of the source $>0.52$ pc (for a screen at 10 kpc) is derived. This limit is consistent with constraints from VLBI observations and low-frequency non-detection.
  • Finally, we discuss the observations under the purview of the best-fit hypernebula model for the PRS associated with FRB20190520B presented in (Sridhar & Metzger, 2022). This best-fit model predicts a rise in flux with time at 1.5 GHz, which does not agree with the observations presented here.
Ongoing monitoring of known PRSs, along with future detections of new PRSs associated with FRBs, will be crucial in further constraining the emission mechanisms and understanding the role of the local environment in the persistent emission.
We thank the staff of the GMRT that made these observations possible. GMRT is run by the National Centre for Radio Astrophysics of the Tata Institute of Fundamental Research. The National Radio Astronomy Observatory is a facility of the National Science Foundation operated under cooperative agreement by Associated Universities, Inc. M.B is a McWilliams fellow and an International Astronomical Union Gruber fellow. M.B. also receives support from the McWilliams seed grant.