### v2

# Continued radio observations of the persistent radio source associated with FRB20190520B provides insights into its origin

Abstract
Follow-up studies of persistent emission from Fast Radio Burst (FRB) sources are critical for understanding their elusive emission mechanisms and the nature of their progenitors. This work presents new observations of the persistent radio source (PRS) associated with FRB20190520B. We observe a gradual decay in the PRS brightness, which is punctuated by periods of brightening and dimming at both 1.5 GHz and 3 GHz. Furthermore, our low-frequency (<1<1 GHz) observations—the first for this source—reveal evidence of a spectral break which can be attributed to absorption processes. Interpreted within the framework of the magnetar wind nebula model, our data constrain the age of the magnetar progenitor to $52^{+16}_{-10}$ years, broadly consistent with previous work. Assuming the observed 1.5 GHz variability is driven by scintillation, we discuss the constraints on the size of the persistent source. The observations presented here challenge the predictions of the previously published best-fit hypernebula model for this source.
software: CASA (2022PASP..134k4501C), CASA CAPTURE (2021ExA....51...95K), emcee (2013PASP..125..306F), ChainConsumer (Hinton2016), Astropy (astropy:2013; astropy:2018; astropy:2022), NumPy (harris2020array), SciPy (2020SciPy-NMeth), Matplotlib (Hunter:2007)
1 Introduction
Fast Radio Bursts (FRBs) are highly energetic $\sim$ millisecond duration radio transients (2007Sci...318.. [^7] L). The observations of FRBs using various radio facilities have revealed that some FRBs repeat, while others appear to be non-repeating (2022A&ARv..30....2P). Among the FRBs that repeat, only two have shown evidence of a periodicity in their activity windows (2022A&ARv..30....2P, and refs. therein). FRBs experience propagation effects due to interaction and scattering before reaching our detectors. Some propagation effects include: dispersion measure (DM)—the delay in arrival of low frequency radiation compared to the high frequency radiation due to interaction with ionized medium; rotation measure (RM)—Faraday rotation of the polarization angle due to magnetic field along the line of sight. The measured DM of the FRB can be expressed as a sum of the DM contribution from the Milky Way, the interstellar medium, the intergalactic medium, and the local DM from the host galaxy of the FRB.
The origin of FRBs is still an open question, with many proposed models (2019PhR...821....1P). The discovery of an FRB-like burst from galactic magnetar SGR1935+2154 has bolstered interest in magnetar-based models as the leading candidates for explaining FRBs (2020Natur.587...59B; 2020Natur.587...54C). To date, a total of $\approx 1000$ FRB sources have been identified, of which $10\%$ are repeaters (2021ApJS..257...59C; 2024ApJ...969..145C; 2023Univ....9..330X; universe11070206). Less than a 100 FRBs have been localized to their host galaxies, providing rich insights into FRB environments (see e.g. 2017Natur.541...58C; 2021ApJ...908L..12T; 2020Natur.577..190M; 2022Natur.606..873N). Among the population of repeating FRBs, a handful of sources are of particular interest given that a compact persistent radio source (PRS) have been observed associated with them—FRB20121102 (2017Natur.541...58C), FRB20190520B (2022Natur.606..873N), FRB20201124A (2024Natur.632.1014B; 2025ApJ...993..234B) and FRB20190417A (2025arXiv250905174M) and FRB20181030A (2024ApJ...976..199I).
FRB20190520B was discovered by the Five-hundred-meter Aperture Spherical radio Telescope (FAST) on May 20, 2019 as part of the CRAFT Survey (2018IMMag..19..112L). It has been localized to a star-forming dwarf galaxy at a redshift of $z=0.241$ (2022Natur.606..873N; 2025ApJ...982..203C). The average DM of bursts from this source is 1207 cm-3pc. Initially, the DM contribution from the host was thought to be exceptionally large $\sim 900$ cm-3pc. However, 2023ApJ...954L...7L investigated the effect of foreground galaxies and estimated a substantially lower host DM. (also see Bhardwaj_2025). Observations of bursts from this FRB using the Green Bank Telescope (GBT) and Parkes telescope by (2023Sci...380..599A) have shown that the RM changes sign twice, pointing to a change in the magnetic field component parallel to the propagation, possibly due to a turbulent magnetized screen surrounding the FRB source. 2023ApJ...958L..19B used Very-long-baseline interferometry (VLBI) observations to constrain the size of the PRS <9<9 pc and co-located it to within $80$ pc of the FRB. This PRS exhibits a flat spectrum (slope of 0.4-0.4 at 2020 and 0.33-0.33 at 2021 epochs) above 1 GHz accompanied by a $\sim 20$% decline in broadband flux between the 2020 and 2021 epochs (2023ApJ...959...89Z). The PRS has not been detected at frequencies <1<1 GHz in the past.
Some leading models proposed to explain persistent radio emission from FRB sources are—Magnetar Wind Nebula (MWN) model (2018ApJ...868L...4M) and hypernebula model (2022ApJ...937....5S). The Magnetar Wind Nebula model attributes persistent radio emission to synchrotron emission from magnetized electron-ion interactions with the nebula, powered by flaring activity of a young magnetar. The Hypernebula model explains the persistent emission as the interaction of winds emitted by an accreting black hole with the surrounding medium. Details of the interpretations of these models are discussed in Section 3.
In this paper, we present the first ever low frequency observations of the persistent radio source associated with FRB20190520B. We interpret the temporal and spectral evolution of the persistent emission by combining our new observations and previous observations of the source. Section 2 describes the observations and data reduction performed, followed by Section 3 covering the analysis of the observations. Finally, we conclude with a discussion of our analysis in Section 4.
2 Observations
2.1 uGMRT observations and archival data
We observed the PRS associated with FRB20190520B using the wideband receiver backend of upgraded Giant Metrewave Radio Telescope (uGMRT) in three frequency bands: band 3 (central frequency, $\nu_{c}$ = 400 MHz, bandwidth, BW = 200 MHz), band 4 ($\nu_{c}$ = 750  MHz, BW = 400 MHz) and band 5 ($\nu_{c}$ = 1260 MHz, BW = 400 MHz) between 2023 June 16 and 2023 June 20 (Proposal 44_039, PI: Balasubramanian). The band 3 observations under this project were split over two days. We also obtained three epochs of observations each in band 4 and band 5 between 2024 September 05 and 2024 September 21 (Proposal 46_126, PI: Balasubramanian). In addition to our observations, we included archival uGMRT data from proposal 43_054 (PI: Yi Feng) in our analysis. Raw data were downloaded in the FITS format and converted to the CASA (2022PASP..134k4501C) measurement set format. The data were then calibrated and imaged using the automated continuum imaging pipeline CASA-CAPTURE (2021ExA....51...95K). All observations used 3C286 as the flux calibrator. J1543-0757 was used as the phase calibrator for observations under 43_054 and 46_126 while J1558-1409 was used as the phase calibrator for observations under 44_039. For the band 3 data, the two measurement sets (under 44_039) were passed onto the automated pipeline separately and then imaged after combining the calibrated measurement sets. Each pipeline run included eight rounds of self-calibration. The band 5 flux density of the persistent source was calculated using the CASA task imfit within a small circular region (of radius $\sim 2\times$ the size of the synthesized beam at band 5) centered at the PRS coordinates. An additional 5% flux density error was added in quadrature to the error obtained from imfit to account for flux calibration errors (see Table 1). No detections were made in Band 3 or in any epoch of Band 4 observations. The upper limit values listed in Table 1 are the $3\timesRMS$ value within a large circular region (of radius $\sim 20\times$ the size of the synthesized beam at the respective band) centered at the position of the PRS. None of the upper limits are used in the analyses in the section that follows. The observation on 2022 Nov 11 (under GMRT 43_053) shows a possible phase calibration issue, which was identified by imaging the flux calibrator using the same calibration solution (see Figure 5 in appendix for a comparison of the non-detection versus detection of the PRS and the flux calibrator). Additionally, the two observations on 2024 Sep 15 and 2024 Sep 21 (under 46_126) were affected by severe radio frequency interference, not allowing for sufficiently clean images.
2.2 VLA observations and archival data
We conducted a single L-band ($\sim$1.5 GHz) observation of the PRS using the Jansky Very Large Array (VLA) on 2024 August 08. In addition, VLA S-band ($\sim$3 GHz) data of the PRS20190520B field were obtained from the archival data listed under proposal 23A-010 (PI: Yi Feng, taken between 2023 June 07 and 2023 June 24). We obtained the calibrated measurement set when available or started with the raw data. Raw data were calibrated using the automated VLA calibration pipeline  [^1] https://science.nrao.edu/facilities/vla/data-processing/pipeline/CIPL_654. All observations used 3C286 as the flux calibrator. J1543-0757 was used as the phase calibrator for observations under 23A-010, while J1558-1409 was used as the phase calibrator for observations under 24A-409. Following calibration, the data were imaged using the automated imaging pipeline  [^2] https://science.nrao.edu/facilities/vla/data-processing/pipeline/vipl_666 and refined through self-calibration. The best self-calibrated image was selected, and the CASA task imfit was used in a small region around the PRS to estimate the source flux density. An additional 5% flux uncertainty was added in quadrature to account for flux calibration errors. It may be noted that the data under 23A-010 has already been analyzed using custom software based on the Astronomical Image Processing System [^3] https://www.aips.nrao.edu/index.shtml by 2024ApJ...976..165Y. Here, we repeated the analysis of this data using the CASA VLA calibration pipeline to maintain consistency with the other measurements in this work. There are slight differences in the obtained measurements compared to the ones discussed in 2024ApJ...976..165Y due to the different calibration and imaging tools used here.
A summary of all measurements is listed in Table 1. Figure 1 shows the decay of the overall flux density of the PRS over time. Figure 2 shows the temporal variation of the PRS flux density at different frequency bands, and Figure 3 displays the variation of the spectrum over the epochs. We observe flaring and dimming episodes of the source in the $\approx 1.5$ GHz data, especially in the late 2020 epoch. The possible reason for this observation is discussed in the Section 3.3. For some observations, the image was heavily affected by radio frequency interference (RFI), resulting in an image that was insufficiently clean to use for our analysis (see Table 1 footnote).
[…]
Table 1: Summary of radio continuum observations of the persistent radio source associated with FRB20190520B. For non-detections, 3 $\sigma$ upper limits are reported, where $\sigma$ is the RMS flux density measured in a large region of the residual image. Observation dates (in two formats), frequency bands, flux densities and their associated errors, and proposal identifiers/paper references are listed. footnotetext: Probable issue with phase calibration. See non-detection in Figure 5.††††footnotetext: Data affected by severe radio frequency interference. Sufficiently clean image was not obtained.
Date Time (MJD) Frequency (GHz) Flux density (uJy) Flux density error (uJy) Reference
2020 Jul 21 59051.067 1.5 258 29 2023ApJ...959...89Z
2020 Jul 23 59053.08 1.5 273 37
2020 Aug 18 59079.019 5.5 145 17
2020 Aug 18 59079.986 5.5 164 19
2020 Aug 29 59090.956 5.5 158 17
2020 Aug 30 59091.953 3.0 195 24
2020 Sep 12 59104.881 3.0 160 21
2020 Sep 12 59104.923 5.5 151 17
2020 Sep 13 59105.991 3.0 186 24
2020 Sep 15 59107.93 5.5 153 17
2020 Sep 19 59111.13 3.0 176 25
2020 Nov 08 59161.691 5.5 139 20
2020 Nov 14 59167.652 3.0 233 29
2020 Nov 16 59169.655 3.0 211 25
2021 Oct 01 59488.879 10.0 115 24 2023ApJ...959...89Z
2021 Oct 01 59488.883 5.5 114 28
2021 Oct 01 59488.887 3.0 112 34
2021 Oct 01 59488.895 1.5 240 70
2021 Nov 07 59525.861 10.0 81 18
2021 Nov 07 59525.865 5.5 139 33
2021 Nov 07 59525.869 3.0 111 33
2021 Nov 07 59525.877 1.5 212 61
2022 Feb 26 59636.208 1.7 197 34 2023ApJ...958L..19B
2022 Feb 27 59637.208 1.7 210 34
[…]
2022 Nov 29 59912.125 1.3 268 44
2022 Dec 27 59940.131 1.3 156 26
2023 Jan 24 59968.036 1.3 281 41
2023 Feb 21 59996.167 1.3 119 36
2023 Mar 21 60024.962 1.3 280 25
2023 Jun 07 60102.323 3.0 157 9 VLA 23A-010 (This work)
2023 Jun 15 60110.234 3.0 173 11
2023 Jun 16 60111.84 0.7 <<165  $\cdots$ GMRT 44_039 (This work)
2023 Jun 17 60112.192 3.0 171 12 VLA 23A-010 (This work)
2023 Jun 18 60113.178 3.0 166 10
2023 Jun 18 60113.219 3.0 165 11
2023 Jun 19 60114.196 0.3 <<195  $\cdots$ GMRT 44_039 (This work)
2023 Jun 20 60115.172 3.0 152 10 VLA 23A-010 (This work)
2023 Jun 20 60115.214 3.0 151 11
2023 Jun 20 60115.255 3.0 143 10
2023 Jun 20 60115.297 3.0 152 10
2023 Jun 20 60115.584 1.3 179 18 GMRT 44_039 (This work)
2023 Jun 23 60118.262 3.0 147 13 VLA 23A-010 (This work)
2023 Jun 24 60119.168 3.0 155 15
2023 Jun 24 60119.209 3.0 144 13
2024 Aug 08 60530.985 1.5 221 13 VLA 24A-409 (This work)
2024 Sep 03 60556.588 1.3 230 24 GMRT 46_126 (This work)
2024 Sep 05 60558.585 0.7 <<155  $\cdots$
2024 Sep 12 60565.589 1.3 184 16
[…]
2024 Sep 21 60574.329 1.3 190 20
[…]
3 Physical Interpretations
Equipped, with the newly observed multi-frequency dataset of the PRS associated with FRB20190520B, we present constraints on the physical properties of the possible progenitor of the persistent emission.
3.1 Monotonic trend in the observations
To investigate presence of a monotonic trend in the light curve, we took all the observations listed in Table 1, and scaled the flux densities to 1.5 GHz assuming a spectral index of 0.4-0.4 (as observed in 2023ApJ...959...89Z, for the 2020 epoch). This scaled data was then subjected to the Mann–Kendall test using the pymannkendall package (Hussain2019pyMannKendall). The test yielded a p-value of $1.9\times 10^{-5}$ and a Kendall Tau value of 0.4-0.4, suggesting a significant decreasing trend. Figure 1 shows the light curve with the scaled flux density values along with the linear regression fit (slope =15.3=-15.3) performed with scipy package (2020SciPy-NMeth) showing the decreasing trend. Changing the assumed spectral index to 0.33-0.33 (as observed in 2023ApJ...959...89Z, for the 2021 epoch) results in similar Mann Kendall p-value ($2.8\times 10^{-4}$), Kenadall Tau value (0.4-0.4) and linear regression fit (slope =11.8=-11.8) as above. This establishes a decreasing trend in the total flux density of the PRS as a function of time.
Figure 1: Light curve with all the flux density measurements from Table 1 scaled to 1.5 GHz assuming a spectral index of -0.4. The fit is a linear regression showing a decreasing trend in the overall flux density over time.
3.2 Magnetar Wind Nebula Model
2016MNRAS.461.1498M and 2018ApJ...868L...4M have shown that a single expanding magnetized electron-ion nebula powered by a young magnetar can interact with the surrounding medium to produce synchrotron emission. The magnetar is assumed to inject energy into the nebula at a rate given by
$\dot{E}=(\alpha-1)\frac{E_{B_{\star}}}{t_{0}}\left(\frac{t}{t_{0}}\right)^{-\alpha}\hskip 5.0pt\rm{for}\hskip 5.0ptt\geqslant t_{0},\hskip 5.0pt\alpha>1$ (1)
where $E_{B_{\star}}$ is the free magnetic energy of the magnetar, $\alpha$ is a powerlaw index, and $t_{0}$ is the onset of the active period of the magnetar. The energy injection drives electrons through the surrounding medium, producing the observed persistent emission. At frequencies $\nu$ above the characteristic synchrotron self-absorption frequency $\nu_{SSA}$, the synchrotron luminosity decays as follows (see 2018ApJ...868L...4M, for detailed calculations)
$L_{\nu}\propto\nu^{-\left(\frac{\alpha-1}{2}\right)}t^{-\frac{\alpha^{2}+7\alpha-2}{4}}$ (2)
We adopt Equation 2 and modify it as follows
$F(\nu,t)=A\,\nu^{-\left(\frac{\alpha-1}{2}\right)}\left(\frac{t_{\rm{obs}}+t_{\rm{age}}}{t_{\rm{age}}}\right)^{-\left(\frac{\alpha^{2}+7\alpha-2}{4}\right)},$ (3)
[…]
Figure 2: Light curve of the FRB20190520B PRS, showing the joint spectro-temporal fit using Equation 3. This fit gives an estimate of the age of the persistent source to be $t_{\rm{age}}=52$ years.
Figure 3: Spectrum of the PRS of FRB20190520B showing the joint spectro-temporal fit using Equation 3. The color of the dashed lines represent the different epochs. The shaded blue region shows the uGMRT upper limits that suggest a break in the spectrum at frequencies <1<1 GHz.
Table 2: MCMC posterior best fit parameters for the decay using Equation 3.
Parameter Prior type Bounds Posterior value
$\alpha$ uniform (1, $\infty$) $1.77^{+0.07}_{-0.08}$
$t_{\rm{age}}$ (yr) uniform (4,1900) $52^{+16}_{-10}$
AA (uJy) uniform (1, $\infty$) $284^{+18}_{-19}$
In addition to the flux decay observed at frequencies $\nu>1$ GHz, the upper limits from our uGMRT observations at $\nu<1$ GHz (from the 2023 and 2024 epochs) suggest a spectral break between 700 MHz and 1.3 GHz (see blue shaded portion in Figure 3). At the observed timescales, such a spectral break is consistent with synchrotron self-absorption. However, the available upper limits are insufficient to constrain the synchrotron self-absorption frequency. Deeper observations at sub-GHz frequencies will be required to quantitatively determine the location of the spectral break. A recent study of the persistent emission from FRB20240114A using GMRT, VLA and MeerKAT data, shows evidence of synchrotron self-absorption peak at $\nu\approx 1.6$ GHz (2025arXiv250114247Z), thereby agreeing with the spectral break below $1.3$ GHz that we observe for the PRS associated with FRB20190520B in this work.
3.3 Scintillation as a source of PRS variability
2023ApJ...959...89Z analysed temporal variability above 1 GHz and found no significant variation in most bands. A marginal variability was seen at 3 GHz. We extend the same variability analysis, adding our 1.3 GHz observations to the 1.5 GHz observations reported by 2023ApJ...959...89Z. One of the causes of variability of a compact radio source is interstellar scintillation. Scintillation arises due to inhomogeneities in the ionized ISM and can cause apparent flux variations in compact radio sources. We follow the formalism introduced in 1992RSPTA.341..151N and further explored in 1998MNRAS.294..307W for our analysis here. The scattering strength parameter $\xi$ describes the scintillation due to a screen at a distance $D$ from the observer. $\xi=1$ corresponds to the critical value at which phase changes are substantial, across the characteristic first Fresnel zone $\theta_{F}=\sqrt{c/2\pi\nu D}$, where $\nu$ is the frequency of observation. Now, $\xi$ is expressed as $(\nu_{0}/\nu)^{17/10}$, where $\nu_{0}$ is the transition frequency between the weak scattering ($\xi\ll 1$ and $\nu_{0}<\nu$), and strong scattering ($\xi\gg 1$ and $\nu_{0}>\nu$) regimes. This transitional frequency has been estimated to be $\nu_{0}=12.53$ GHz using the galactic free electron density model in pyNE2001 (2002astro.ph..7156C). Given $\nu_{0}=12.53$ GHz, we are in the strong scattering regime as $\nu_{0}>\nu$ (as $\nu=1.5$ GHz). There are two possible scintillation scenarios: diffractive scintillation and refractive scintillation. Diffractive scintillation is relevant at characteristic frequency bandwidth, of $\xi^{-2}\nu=\nu(\nu_{0}/\nu)^{17/5}$ which is $\sim 1$ MHz for our observations. Our observations span a larger bandwidth, washing out diffractive scintillation. Hence, refractive scintillation is the relevant mechanism here. For, refractive scintillation, the expected modulation index is
$m_{\rm{exp}}=\xi^{-1/3}=\left(\frac{\nu}{\nu_{0}}\right)^{17/30}=0.29$ (4)
Assuming a scattering disk at a distance of $D=1$ kpc from us, the size of the Fresnel zone at $\nu=1.4$ GHz is $\theta_{F}=\sqrt{c/2\pi\nu D}=6.78\,\mu\rm{as}$. Therefore, the angular size of the screen is
$\theta_{r}=\theta_{F}\xi=284.1\,\mu\rm{as}$ (5)
This angular size of the screen can be converted to a projected radius of the screen at the distance of the source using the luminosity distance of the host galaxy (i.e. $d_{\rm{lum}}=1218$ Mpc) as follows:
$R_{r}=\theta_{r}d_{\rm{lum}}/2=5.0\times 10^{18}\,\rm{cm}=1.68\,\rm{pc}$ (6)
The characteristic timescale associated with the refractive scintillation from this screen is
$t_{r}=2\left(\frac{\nu_{0}}{\nu}\right)^{11/5}=10.47\,\rm{days}$ (7)
Given this expected timescale, we compute the modulation index of the observed data in the frequency range 1.3–1.7 GHz from Table 1, by averaging consecutive data that are within the refractive scintillation timescale, $t_{r}$, using the following expression:
$m_{\rm{obs}}=\frac{1}{\overline{F_{i}}}\sqrt{\frac{N}{N-1}\left(\overline{F_{i}^{2}}-\overline{F_{i}}^{2}\right)}$ (8)
where $F_{i}$ are the flux density measurements, $\overline{F_{i}}$ is the mean of the flux density measurements, $\overline{F_{i}^{2}}$ is the mean of the squares of $F_{i}$ and $N$ is the total number of observations.
We obtain the observed modulation index $\bm{m_{\rm{obs}}=0.23}$. The observed modulation of the PRS flux is less than that expected from this region of the sky, i.e. $\bm{m_{\rm{obs}}<m_{\rm{exp}}}$. This observation can be interpreted in the following ways:
  • All the variability observed is due to scintillation, and the pyNE2001 estimate of the transition frequency is accurate. However, the observed flux modulation is lesser than the expected modulation. This change in modulation can be explained if the size of the source $\theta_{s}$ is slightly larger than the size of the screen $\theta_{r}$, reducing the modulation index by a factor of $(\theta_{r}/\theta_{s})^{7/6}$, and increasing refractive timescale by a factor of $\theta_{s}/\theta_{r}$ (see 1998MNRAS.294..307W, for details). Table 3 summarizes the possible constraints on the size of the source ($R_{s}=\theta_{s}d_{\rm{lum}}/2$), and the modified timescale of the scintillation, for screens at a distance of 0.1 kpc, 1 kpc and 10 kpc. The projected screen radius ($R_{r}$) values listed in Table 3 can be quoted as a conservative lower limit of the size of the source. These are shown in Figure 4 for different distances of the screen from the observer in the $t_{\rm{age}}$$R_{n}$ phase space plot, adapted from 2023ApJ...958L..19B.
  • All the variability observed is due to scintillation, but the pyNE2001 model estimate of the transition frequency is not accurate. If this is the case, we can estimate the transition frequency required to produce the observed modulation. Table 3 lists the expected transition frequency, $\nu_{0}$ which can produce the observed modulation index at $\nu=1.4\,$GHz.
  • All the variability observed is intrinsic to the source. This implies that the size of the source $\theta_{s}$ is larger than the size of the screen $\theta_{r}$ and therefore, refractive scintillation (which assumes a point source) is not relevant here. The projected size of the screen $R_{r}$ tabulated in Table 3 can be used as a conservative lower limit of the size of the source in this case.
At this point, we also note that there is a possibility of both intrinsic variability and scintillation to contribute to the observed variability. It is not trivial to separate the two without independent constraints on the size of the source and the properties of the scattering screen.
For completeness, we also repeat this analysis after removing the expected decay from the magnetar wind nebula model (Section 3.2). The inferred size and timescale constraints remain broadly consistent in both cases.
Figure 4: Constraint on the radius of the PRS source using the projected size of the screen $R_{r}$. The figure is adapted from 2023ApJ...958L..19B, after correcting for the radius by dividing all limits by 2.
Table 3: Scintillation analysis and predictions at $\approx$1.4 GHz (the observation frequency used is a mean of all the detections from Table 1 i.e. $\nu\approx 1.4$ GHz)
Dataset $t_{r}$ (days) $m_{\rm{exp}}$ $m_{\rm{obs}}$ Characteristic sizes modified $t_{r}$ [1em] (days) modified $\nu_{0}$ [1em] ( GHz)
$D=0.1$ kpc $D=1$ kpc $D=10$ kpc
[…]
Data as it is 10.47 0.29 0.23 21.44 5.30 6.48 6.78 1.68 2.05 2.14 0.53 0.65 12.80 18.93
Removed best fit decay 0.22 6.80 2.15 0.68 13.43 20.92
3.4 Hypernebula Model
Another model proposed to explain the PRSs associated with FRBs is the hypernebula model (2022ApJ...937....5S). In this model, mass transfer from a companion star onto a massive black hole via Roche lobe overflow drives powerful disk winds. These winds inflate a bubble of relativistic electrons that emit synchrotron radiation. Two key timescales characterize the model: the timescale over which the wind expands freely into the CSM, $t_{\rm{free}}$; and $t_{\rm{active}}\sim 10$$10^{6}$ years, the total duration of the accretion phase (2022ApJ...937....5S). The relatively modest evolution in burst DM over time (2022Natur.606..873N; 2023Sci...380..599A) suggests that the system is currently at $t<t_{\rm{free}}$. However, 2021ApJ...923L..17Z show a decreasing trend in the burst DM. (2022ApJ...937....5S) applied the hypernebula model specifically to FRB20190520B, using a set of best-fit model parameters (e.g., age, mass accretion rate, ambient density) chosen to match the source’s known properties at the time. For those specific parameters, their model predicts a gradual increase in radio flux of the PRS at GHz frequencies over decadal timescales—this is explicitly stated in Section 4.2 of their paper and shown in their model curves (see Figures 5 and 6 in 2022ApJ...937....5S). Whereas, we see a slow decay in the flux over the span of $\sim 4$ years. This is in contrast to our observations, indicating that the best-fit hypernebula model using data from the first 2 years since detection is inconsistent with the recent observed behavior of the PRS associated with FRB20190520B.
3.5 AGN or IMBH Origin for the Persistent Radio Source
An alternative scenario is that the persistent radio source (PRS) associated with FRB20190520B is physically unrelated to the FRB engine, and instead arises from accretion onto a massive black hole as discussed by 2023Sci...380..599A. In this picture, the PRS represents compact synchrotron emission from a radio-loud, low-Eddington active galactic nucleus (AGN), while the FRB originates from a separate compact object—such as a magnetar—embedded within the same nuclear environment.
Recent work by 2024ApJ...973..133D has systematically compared the radio spectra of persistent sources in nearby dwarf galaxies—including known low-mass AGNs—with those of FRB-associated PRSs, including FRB20190520B and FRB 20121102A. They find that several AGNs in dwarf hosts exhibit flat or slightly inverted radio spectra and luminosities consistent with both FRB PRSs, suggesting that an AGN origin cannot be ruled out based on spectral properties alone.
[…]
A similar interpretation has been proposed for the PRS associated with FRB 20121102A. 2025PASP..137h4202B argued that its long-term flux stability, compact size (<0.35<0.35 pc), and placement on the radio-loud fundamental plane of black hole activity can be explained by a low-Eddington AGN powered by an IMBH. Given the morphological and spectral similarities between the two sources, a comparable AGN origin for the PRS of FRB20190520B remains observationally viable. However, as with FRB 20121102A, this model requires that the FRB bursts themselves arise from a distinct, co-located source within the broader accretion environment.
Further high-resolution radio monitoring and deep X-ray observations will be essential to distinguish between AGN and magnetar wind nebula interpretations and to assess whether FRB sources and persistent emitters are causally connected or merely cohabit similar extreme environments.
4 Conclusion
In this paper, we present broadband radio observations (0.30.3–-$10$ GHz) of the persistent radio source (PRS) associated with FRB20190520B. We analyze the data under different proposed models and comment on the physics responsible for the observed emission. The findings are summarized below.
  • We observe a slow decay in the brightness of the PRS with time across all frequency bands. There are episodes of brightening and dimming seen in the 1.5 GHz and 3 GHz data.
  • The first low frequency (<1<1\,GHz) observations of this PRS hint at a spectral break between $\approx 0.7$ GHz and $\approx 1.3$ GHz, which can be attributed to synchrotron self-absorption. It is not possible to estimate the self-absorption frequency using only the upper limits from the low-frequency observations presented here.
  • We fit the decaying light curves at all frequencies using the decay equation described in Equation 3, derived from the magnetar wind nebula model. This resulted in an estimate of the powerlaw index for energy injection $\alpha=1.77^{+0.07}_{-0.08}$ and the age of the magnetar is $t_{\rm{age}}$, is $52^{+16}_{-10}$ years. These values are in agreement with constraints from previous observations of the PRS (2023ApJ...958L..19B; 2021ApJ...923L..17Z; 2024arXiv241219358B).
  • We also comment on the contribution of scintillation to the variability of the PRS at 1.5 GHz. Through this analysis, a conservative lower limit on the size of the source >0.53>0.53 pc (for a screen at 10 kpc) is derived (assuming that NE2001 accurately models the transition frequency and that the observed variability is due to scintillation). This limit is consistent with constraints from VLBI observations and low-frequency non-detection. We also note that the variability could be a combination of scintillation and intrinsic variability, and it is not trivial to distinguish between the two.
  • Finally, we discuss the observations under the purview of the best-fit hypernebula model for the PRS associated with FRB20190520B presented in (2022ApJ...937....5S). This best-fit model predicts a rise in flux with time at 1.5 GHz, which does not agree with the observations presented here.
Ongoing monitoring of known PRSs, along with future detections of new PRSs associated with FRBs, will be crucial in further constraining the emission mechanisms and understanding the role of the local environment in the persistent emission.
[…]
Appendix A Comparison of calibration of images
This section contains a comparison of the image of the target PRS along with the calibrator, 3C286 to demonstrate the calibrations errors present in the 2022 November 11 data. This is shown in Figure 5.
Figure 5: Comparison of the image of the target PRS (top row) and the calibrator 3C286 (bottom row) on 11 Nov 2022 (left column) and 29 Nov 2022 (right column). This shows the calibration errors in the 11 Nov 2022 data.